A 59-m rope is cut into three pieces. The second piece is twice as long as the first. The third piece is 3m long than the first. How long is each piece of rope?
Answer by scott8148(5895)

let x=first piece

"The second piece is twice as long as the first" __ 2x

"The third piece is 3m long than the first" __ x+3

"A 59-m rope is cut into three pieces" __ 59=x+2x+x+3 __ 59=4x+3 __ 56=4x __ 14=x